Solve system of 3 equations:
15x-25y+z+10=0
30x-50y+2z+20=0
90x-150y+6z+60=0
Can you show steps by steps

Answer by Alan3354(69443)

Solve system of 3 equations:
15x-25y+z+10=0
30x-50y+2z+20=0
90x-150y+6z+60=0
Can you show steps by steps
================
The 2nd equation is 2 times the 1st, the 3rd is 6x the 1st
--> dependent, no unique solution.